4.5 mL of 0.0750 M KOH is added to a 32.0 mL solution of 0.025 M HNO2. Calculate the pH of the solution.
Answer #1

KOH + HNO2 ------> KNO2 + H2O

moles of KOH = molrity*volume of solution in litres = 0.075*0.0045 = 3.375*10-4

moles of HNO2 = molrity*volume of solution in litres = 0.025*0.032 = 8*10-4

Thus, excess HNO2 = (8*10-4) - (3.375*10-4) = 4.625*10-4

Now,

HNO2(aq) -------> H+(aq) + NO2-(aq)

[H+] = 4.625*10-4 M

pH = -log[H+] = 3.335